# Learning Loss Landscapes in Preference Optimization

### v1

@@ 3.1 Learning mirror maps
To search for the best $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy (Salimans et al., 2017). Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient.
4 Experiments
We carry out our experiments on continuous reinforcement learning tasks in MuJoCo and on LLM fine-tuning. In particular, due to computational efficiency constraints, we discover mirror maps in MuJoCo and report their performance when evaluated for an LLM finetuning task.
4.1 Environments
We consider three different preference optimization settings which we describe below. Each setting includes a starting policy, a set of preference datasets with different properties and an evaluation method for the trained policy. We apply ES to the first two settings and test the discovered objectives on the last one.
Hopper
Our objective in Hopper is to learn a policy from scratch, i.e. with a randomly initialised policy, using a preference dataset. On Hopper, we train four reference agents of different skill levels, specifically, of respective expected cumulative reward (or value) of 900, 1200, 1800, and 2100 (the expert agent). Each dataset consists of 5120 rows, each with two trajectories of length 1000 starting from the same state. We generate a dataset by comparing trajectories by the expert agent (average return 2100), with other trajectories by one of our reference agents. A Bradley-Terry judge ranks each pair of trajectories, based on their true reward. The trained policy is then evaluated on the standard Hopper environment. To test the efficacy and robustness of our loss functions, we analyse a few variations of these standard datasets that represent common issues of real world data.
  • Base dataset: we compare a trajectory from the expert agent with one of a reference agent.
  • Noisy dataset: we flip a varying portion of the rankings in the base dataset.
  • Shuffled dataset: the comparisons are not always between the expert agent and a second agent. Namely, $25\%$ of the comparisons are between two trajectories from the expert agent, $50\%$ of the comparisons are between a trajectory of the expert agent and one of a reference agent, and $25\%$ of the comparisons are between two trajectories of a reference agent.
  • Poor judge dataset: the judge is more likely to flip labels when the two trajectories have closer reward values. This is implemented as an increase in the temperature of the Bradley-Terry judge.
Ant
To further investigate a scenario similar to fine-tuning a large language model (LLM), we examine a setting in which a pre-trained agent is required to modify its behavior to achieve its goal with a stylistic constraint. Specifically, in the Ant environment, we take an agent that has been pre-trained on the standard Ant goal of moving forward, and enforce the objective of avoiding the use of one of its legs. This is accomplished by introducing the Three-legged-ant (TLA) environment, a modified version of Ant where utilizing the fourth leg results in significant penalties. We train one agent in the original Ant environment, achieving a reward of 6000, and another in the TLA environment, which achieves a reward of 3900. For comparison, the agent trained in the original Ant environment achieves a reward of 1700 when evaluated in the TLA environment. The dataset generation follows the same protocol as described for the Hopper environment, where trajectories are collected from both the Ant and TLA agents. However, in this setting the number of rows for each dataset is 1280, to account for the fact that we do not want the agent to learn from scratch, but to adjust its policy to a new instruction. The trained policy is then evaluated on the TLA environment.
LLM tuning
Finally, we evaluate one of our discovered objective functions on a real-world LLM fine-tuning task. To simulate a scenario involving mixed data quality, we fine-tune the gemma-7b [^2] https://huggingface.co/google/gemma-7b model on a modified version of the dpo-mix-7k [^3] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset, where half of the responses, selected at random, are replaced with responses generated by gemma-2b [^4] https://huggingface.co/google/gemma-2b, a model that typically produces lower-quality responses compared to those originally present in DPO-mix. This approach replicates the shuffled dataset described for the MuJoCo experiments, aiming to simulate the challenges of training with datasets of varying quality, a common issue in fine-tuning tasks where it is difficult and costly to ensure uniformly high-quality data.
We implement our experiments in MuJoCo using the brax library (Freeman et al., 2021) for the environments and the evosax library (Lange, 2022) for the evolution. We use the same hyper-parameters for all MuJoCo datasets, which have been tuned for ORPO using Weights and Biases [^5] https://wandb.ai/site. We report them in Appendix C, along the hyper-parameters for OpenES. For the LLM tuning, we modify the Alignment Handbook library (Tunstall et al., ), and use the default hyper-parameters.
4.2 Results: Mujoco
Table 1: Results for Hopper. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset. When ranking two trajectories of value 2100 and 1800, a judge with low, medium, and high temperature provides the correct ranking 95%, 85% and 75% of the time, respectively.
Value agent 1 Value agent 2 Noise Shuffled Judge temp. ORPO Discovered
2100 900 0 No Low 2003$\pm$20 2012$\pm$11
2100 1200 0 No Low 2055$\pm$12 2055$\pm$10
2100 1800 0 No Low 2043$\pm$ 15 2073$\pm$12
2100 2100 0 No Low 2070$\pm$12 2098$\pm$11
2100 900 0.1 No Low 1519$\pm$45 1724$\pm$43
2100 1800 0.1 No Low 1936$\pm$25 1945$\pm$25
2100 2100 0.1 No Low 2112$\pm$9 2094$\pm$13
2100 900 0.2 No Low 662$\pm$49 1127$\pm$45
2100 900 0.3 No Low 623$\pm$29 698$\pm$26
2100 1800 0 No Medium 1884$\pm$22 1928$\pm$23
2100 1800 0 No High 1858$\pm$32 1895$\pm$14
2100 900 0 Yes Low 975$\pm$51 1099$\pm$35
Table 2: Results for TLA. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset.
Value agent 1 Value agent 2 Noise Shuffled ORPO Discovered
3900 1700 0 No 3277$\pm$49 3473$\pm$63
3900 1700 0 Yes 2425$\pm$58 3485$\pm$163
3900 1700 0.1 No 2675$\pm$62 3837$\pm$52
Table 3: Results for temporally-aware objectives. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the temporally-aware objective discovered for that dataset.
Value agent 1 Value agent 2 Env Noise Shuffled ORPO Discovered
2100 900 Hopper 0 Yes 975$\pm$51 2032$\pm$13
3900 1700 TLA 0 No 3277$\pm$49 3903$\pm$42
3900 1700 TLA 0 Yes 2425$\pm$58 3251$\pm$83
We provide the results of our experiments for MuJoCo in Tables 12, and 3, which report the performance of ORPO and our discovered objectives for Hopper, TLA, and the temporally aware case, respectively. To provide information on the behavior of these objectives as well as guidelines on how to design loss landscapes for the settings we have considered, we compare the landscapes of ORPO and of the discovered objectives in Figures 123, and 4. In these figures, we report, in absolute value, the gradients of the objective with respect to $\log(\pi(\tau_{w}))$ and $\log(\pi(\tau_{l}))$, as well as a sampled training trajectory. We discuss the results below, by dataset type.
Base Dataset
For the base dataset, we observe different behaviors for the Hopper and TLA environments. For the Hopper dataset, we see that ORPO achieves the performance of the best reference agent in all cases and that it is not significantly outperformed by the learned objective. On the other hand, ORPO does not reach the performance of the best reference agent in the TLA dataset and is surpassed by the discovered objectives. Remarkably, the temporally aware objective achieves the performance of the best reference agent, presenting a $19\%$ improvement over ORPO. As shown by Figures 1 and  4, both the static and the temporally aware discovered objectives present a smaller gradient than ORPO, which means that they cause smaller updates. Additionally, for ORPO we have that if $\log(\pi(\tau_{w}))<\log(\pi(\tau_{l}))$ then the gradients increases suddenly by half, while for the temporally aware discovered objective we can observe that the gradient increases more slowly. Essentially, this means that ORPO causes a large update when the agent mistakenly assigns more probability to the rejected trajectory, while the temporally aware discovered objective causes a large update only if the agent assigns significantly more probability to the rejected trajectory.
Figure 1: The plots represent the gradients, in absolute value, of ORPO (left) and the objective discovered on the base TLA dataset (right). The dots represent a sampled training trajectory.
Noisy Dataset
As expected, the performance of ORPO declines when the dataset contains randomly flipped labels, with the degradation becoming more pronounced as the percentage of flipped labels increases. While the performance degrades for our discovered objectives in Hopper as well, it happens more slowly and we show significant improvements over ORPO for all noise levels. In TLA, the discovered objective is capable of almost achieving the best reference agent value despite the noise. Figure 2 shows that the objective discovered in Hopper assigns a roughly constant value to the gradient w.r.t. the chosen logits, while the gradient w.r.t. the rejected logits increases as $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ decreases. Consequently, when the agent encounters a flipped label where $\log(\pi(\tau_{w}))\ll\log(\pi(\tau_{l}))$, the agent does not take larger updates to increase the chosen logits, but only larger updates to decrease the rejected logit, thus mitigating the noise effects. On the other hand, the objective discovered in TLA manages to mitigate the influence of noise by reducing the gradient value and by taking larger updates only when the chosen logits are much larger than the rejected logits.
Figure 2: The plots represent the gradients, in absolute value, of the objectives discovered on the noisy Hopper dataset (left) and on the noisy TLA dataset (right). The level of noise is 0.1 for both plots. The dots represent a sampled training trajectory.
Shuffled Dataset
While effectively using the same data as the base dataset, the performance of ORPO significantly decreases in both the Hopper and TLA shuffled dataset, where low quality trajectories are selected as the preferred choice when compared to other low quality trajectories. In contrast, the discovered objectives achieve a value of the final policy close to the best reference agent. In particular, the value reached by the temporally aware objective for Hopper is more than double the value reached by ORPO. All discovered objectives present smaller gradients than ORPO, and therefore induce smaller updates, which prevents taking large updates when the chosen logits belong to a low-quality trajectory. As to the temporally aware objective discovered on HopperFigure 4 shows that it induces very small updates for the chosen logits when they are low, and therefore probably belong to a low-quality trajectory. Lastly, we note a decrease in the absolute value of the gradient at the top and right borders as training progresses, which induce smaller updates.
Figure 3: The plots represent the gradients, in absolute value, of the objectives discovered on the shuffled Hopper dataset (left) and on the shuffled TLA dataset (right). The dots represent a sampled training trajectory.
Poor judge
As expected, we observe a decrease in performance for ORPO when we increase the temperature of the judge. However, since we compared trajectories from two high-quality reference agents, the drop in performance is small. This is also reflected in the discovered objectives, which do not manage to significantly outperform ORPO.
Figure 4: The plots represent the gradients, in absolute value, of the temporally aware objectives discovered on the shuffled Hopper dataset (left) and on the base TLA dataset (right), for three different values of the training progress. For each row, the dots represent the respective third of a sampled training trajectory.
4.3 Results: Transfer
We show that PO objectives learnt with our methodology can transfer to other domains. In particular, the static objective discovered on the shuffled TLA dataset tranfers to both the shuffled Hopper dataset, where it obtains a final policy value of 1735$\pm$39, and to the LLM-tuning task with shuffled data defined above. We train the base LLM on the modified DPO-mix dataset with both ORPO and the discovered objective. We observe that the first methodology achieves 57% accuracy on the test set while the latter achieves 62%. Additionally, we compare the two trained models with AlpacaEval (Li et al., 2023), and obtain a 53% winrate for the model trained with the discovered objective.
5 Related Work

### v2

@@ 3.1 Meta Learning PO objectives
To search for the best $g$, $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy. Denote by $\zeta$ the parameters of $g$, $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient. In practice, to compute (12), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions. We then train 256 agents with the perturbed objective functions on an a preference dataset. To measure the value of each agent, i.e. $\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent. Refer to Appendix E for further discussion and details on the ES methodology.
In our experiments, we consider both the case where we fix $g=\log$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions. We also perform the evolution on both the two-step and the one-step MPO classes.
Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
RRHF 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
SLiC-HF 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
DPO 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
IPO 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
CPO 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
ORPO 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
R-DPO 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
SimPO 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
With $g=\log$
1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
With parametrized $g$
1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
Table 2: Three Legged Ant (TLA) Adaptation. Performance of popular and discovered MPO algorithms on TLA, for various dataset settings. For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents. For discovered MPO algorithms, we report performance of one-step MPO (1S-MPO) and two-step MPO (2S-MPO). For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters). We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
4 MuJoCo Experiments
Our first set of experiments is carried out on continuous RL tasks in MuJoCo. In particular, we show the performance of all the algorithms presented in Table 1 across several settings and we compare it with the performance of our discovered objectives. To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
4.1 Tasks
To reproduce the typical conditions of LLM fine-tuning, which involve a pre-trained model, we consider a setting where the task is to adapt a pre-trained agent to meet the original objective while adhering to an additional stylistic constraint. Specifically, in the Ant environment, we start from an agent that has been pre-trained on the standard Ant goal of moving forward and enforce the objective of avoiding the use of one of its legs. This is accomplished by introducing the Three-legged-ant (TLA) environment, a modified version of Ant where utilizing the fourth leg results in significant penalties.
The offline preference optimization task is defined as follows. We train one agent (the original agent) in the original Ant environment, achieving a reward of 6000, and another (the target agent) in the TLA environment, which achieves a reward of 3900. For comparison, the original agent achieves a reward of 1700 in the TLA environment. We then generate a preference dataset of 1280 rows, each with two trajectories of length 1000 starting from the same state. Each trajectory is generated by either the original or the target agent, depending on the current setting. A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward.
We also consider training a randomly initialized agent on the Hopper environment, using a preference dataset with 5120 pairs of trajectories. This setting addresses the case where a new behavior has to be learned completely from the preference dataset and no prior knowledge of the task is available. Due to space constraints, we report the results of the experiments for this task in Appendix H.
We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
  • Base dataset: for each pair of trajectories, one is generated by the original agent and one by the target agent.
  • Noisy dataset: same as the base dataset but each label given by the judge is flipped with probability $\varepsilon$.
  • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one. The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between a trajectory of the target agent and one of the original agent, and $25\%$ comparisons between two trajectories of the original agent.
Figure 1: SimPO
Figure 2: Discovered 1S-MPO (shuffled)
Figure 3: ORPO
Figure 4: Discovered 1S-MPO (noisy)
Figure 5: Plots of the absolute value of the gradient of SimPO, ORPO, and of two discovered objectives, i.e. 1S-MPO (shuffled) and 1S-MPO (noisy, $\epsilon=0.3$). The dots are sampled datapoints from the training distribution, with lighter points representing more recent datapoints during the course of training. In the bottom right and corner of each plot, SimPO’s and ORPO’s gradients go to zero, while the gradient of our discovered loss functions keep increasing. The effect of this on training is evident observing the datapoints: in ORPO and SimPO, the likelihood of the rejected points is not minimized, leading to decreased performance.
4.2 Results
We provide the results of our experiments for TLA in  Table 2, which reports the performance of several PO algorithm and of our discovered objectives. We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters. All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
First of all, we notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a large drop in performance in the mixed-quality and noisy settings.
Importance of SFT
The first group within Table 2 shows that SFT plays a key role in the performance across different settings. SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings. IPO and DPO, which do not contain an SFT term but have an SFT step, are among the top performers on the base, mixed-quality and low noise settings. Their performance finally drops when the noise level reaches 0.3. Lastly, the algorithms that present an SFT term in their objectives, e.g. CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
Keep optimizing
Furthermore, while RRHF and SliC-HF are very similar, SliC-HF allows the PO part of the objective to be clipped when $\pi(\tau_{w})>\pi(\tau_{l})+\delta$, rather than when $\pi(\tau_{w})>\pi(\tau_{l})$. This modification leads to a higher performance on all tasks, demonstrating that it is beneficial to keep optimizing the policy even if $\pi(\tau_{w})>\pi(\tau_{l})$. To further stress this point, we consider the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big{]},$
which we call Linear Preference Optimization (LPO). LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table 2, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
Discovered objectives
Table 2 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it. We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings.
Differently from the human-designed algorithms, the discovered objectives recover the performance of the target agent in multiple instances and are more robust to the mixed-quality and noisy settings. We note that allowing the evolution procedure to learn $g$ leads to a better performance in all dataset settings. Additionally, we have that the objectives discovered within the two-step MPO class always have a lower performance than those within the one-step MPO class. This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
When exploring the one-step MPO class with $g=\log$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\alpha\log(\pi(\tau_{w}))$
$+\lambda\log\sigma\left(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\right)\big{]},$
where the coefficients $\alpha$ and $\beta$ depend on the setting. In particular, we have a low value for $\alpha$ and a high value for $\beta$ in the noisy settings, while we observe the opposite in the mixed-quality setting. These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories. When we search the two-step MPO class with $g=\log$, we recover DPO.
Figure 5 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$. In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings. For comparison, we provide the same plots for the ORPO and SimPO objectives. The hand-crafted algorithms present a larger gradient when $\pi(\tau_{w})<\pi(\tau_{l})$ and a smaller one when $\pi(\tau_{w})>\pi(\tau_{l})$, that is, they induce large updates when the data-point contradicts the current behaviour of the agent, and small otherwise.
The objective discovered on the noisy dataset has the opposite behavior, meaning that it only optimizes the more robust-to-noise SFT term when $\pi(\tau_{w})<\pi(\tau_{l})$. As $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ becomes larger, it shifts toward increasingly large updates thanks to the PO term. A similar pattern appears in the mixed-quality dataset, where the objective also increases its updates as the difference $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ grows. The key distinction between these two losses is that the shuffled loss triggers high-gradient updates even when $\pi(\tau_{w})<\pi(\tau_{l})$. We highlight once more that both discovered objectives advocate to keep optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
For additional results and analysis for the Hopper task, we refer the reader to Appendix H.
WR $s_{\text{WR}}$ LCWR $s_{\text{LCWR}}$ avg length
SLiC-HF 16.77 1.32 13.18 0.28 3016
ORPO 24.47 1.52 19.25 0.29 1679
CPO 25.96 1.54 21.07 0.25 1623
CoLPO 31.94 1.65 24.01 0.21 2391
Table 3: Alpaca Eval LLM results. We report Win Rate (WR), standard error for Win Rate ($s_{\text{WR}}$), Length Controlled Win Rate (LCWR), standard error for LCWR ($s_{\text{LCWR}}$) and average length of responses. We trained gemma-7b on dpo-mix-7k for 3 epochs. We then compared our model to GPT-4 using Llama-3.3-70B-Instruct as judge. All algorithms use the same $\beta=0.5$ and the same learning rate of $55e-$ 5. In the case of CoLPO and SLiC-HF, we used the same cutoff rate at 5.
5 Experiments: LLM transfer
We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting. In particular, we found that all our discovered loss functions kept optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, while all algorithms in Table 1 have a PO term whose gradient quickly reaches 0 when $\pi(\tau_{w})>\pi(\tau_{l})$.
Using this insight, we design a new PO objective, which we call Constrained Linear Preference Optimization (CoLPO) and define as
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(\tau_{w}|x)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)<\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),0)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)>\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),\gamma)\big{]},$
where $\beta\geq 0$, $\gamma\geq 0$ and $\alpha\leq 0$ are hyperparameters. The CoLPO objective keeps optimizing the policy when $\pi(\tau_{w})>\pi(\tau_{l})$, but only if $\log\pi_{\theta}(\tau_{w}|x)$ is over a certain threshold. This constraint is placed to stop the optimizer from increasing the difference $\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))$ by descreasing both probabilities.
We evaluate CoLPO on an LLM fine-tuning task, where we fine-tune the gemma-7b (Team et al., 2024) model on the dpo-mix-7k [^2] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset for 3 epochs. Being a real-world preference dataset, dpo-mix-7k naturally presents characteristics of the mixed-quality and the noisy settings. To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the CoLPO objective. We evaluate the tuned LLM against GPT-4, using the AlpacaEval library  (Li et al., 2023). All models were trained using the default hyper-parameters, with $\beta=0.5$, $\gamma=5$ (for CoLPO and SLiC-HF), and $\alpha=-1$ (for CoLPO). For details on hyperparameters and training settings consult Appendix J.
Table 3 reports the results of CoLPO, as well as the results for other one-step algorithms, and shows that CoLPO largely outperforms existing algorithms, confirming the validity of the insights gained in the MuJoCo simulations.
Lastly, we report that the the log-probabilities for the final training batch of chosen and rejected responses are -1.17 and -11.38 respectively for CoLPO, compared to -0.61 and -2.98 for ORPO, respectively.
This results further confirms that CoLPO is effective in decreasing the probability of rejected responses and mantaining high the probability of chosen responses.
6 Conclusion
We have introduced a novel framework for Preference Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolutionary strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we have demonstrated that the performance of our discovered objectives consistently exceeds the performance of existing methods, particularly in noisy and mixed-quality datasets where many baselines underperform.
Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$. To address this, we proposed a simple objective, CoLPO, that avoids such loss truncation while preventing the optimizer from decreasing the probability of chosen datapoints. We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
Appendix A Popular PO Algorithms
These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT.
The comparison also includes IPO, a theoretically-based approach that handles pairwise preferences differently than DPO. Another method, CPO, combines sequence likelihood as a reward with an SFT objective. Finally, R-DPO modifies the original DPO by adding regularization to prevent length exploitation.
Appendix B Related Work